PQ = 3x + 7, QR = 5x − 19


1
Expert's answer
2021-10-15T11:27:02-0400

Notice that, if Q is the midpoint of PR, that simply means that PQ=QR


"PQ=3x+7"

"QR = 5x \u2212 19"

"3x+7=5x-19"

"5x-3x=7+19"

 "2x=26"

"x=13"